OpenStudy (anonymous):

Find the exact value of cos(-7pi/4 radians) (sqrt 2)/(2) 1/2 (sqrt 3)/(2) -1/2

OpenStudy (anonymous):

@jim_thompson5910 @mathstudent55 @Hero

OpenStudy (anonymous):

\[\cos \alpha is periodic with 2 pi\] \[\cos (-7\pi/4)=\cos \left( \frac{ -7\pi }{ 4}+2\pi \right) =\cos \frac{ \pi }{4 }=\frac{ 1 }{ \sqrt{2} } \]

OpenStudy (anonymous):

\[\frac{ 1 }{\sqrt{2} }=\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

Sweet, thanks!